OpenStudy (anonymous):

PLZ HELP how to value of the expression 8^–2

OpenStudy (texaschic101):

a^-2 = 1/a^2

OpenStudy (texaschic101):

knowing this....what is 8^-2 ??

OpenStudy (anonymous):

4

OpenStudy (texaschic101):

8^-2 = 1/8^2 = 1/64.....understand ??

OpenStudy (anonymous):

YES THANKS

OpenStudy (texaschic101):

happy to help...although I happen to think negative exponents are a pain in the but...lol

OpenStudy (madhu.mukherjee.946):

@ajkishon09 dude why don't you people give medals to those who help you out?????????????

OpenStudy (anonymous):

8^-2 = 1/8^2 = 1/64 @texaschic101 is right